# Mitigating photon loss in linear optical quantum circuits: classical postprocessing methods outperforming postselection

### v2

@@ 8.1 Linear solving
1  Initialise variable $p_{\text{miti}}(s^{n}_{l})\leftarrow 0$ for the mitigated output to be computed.
2 Update mitigated output variable as $p_{\text{miti}}(s^{n}_{l})\leftarrow{{m-n+k\choose k}}\big{|}\tilde{p}_{R}^{k}(s^{n}_{l})-\frac{N^{\prime}_{k}}{N_{k}}p_{unif}\big{|}$.
@@
There exists a regime of recycling mitigation usefulness, where the combined bias and the statistical errors present in the mitigated probabilities are lower than the statistical errors of the postselected distribution. By combining upper bounds for each of these errors we will now provide conditions for recycling mitigation using linear solving to outperform postselection.
These are stated in the form
$N_{k}\bigg{(}\frac{N^{\prime}_{k}}{N_{k}}\mathsf{M}\big{(}\epsilon_{bias,s^{n}_{l}}\big{)}+\mathsf{M}\big{(}\epsilon_{stat.,s^{n}_{l}}\big{)}\bigg{)}\leq\mathsf{M}\big{(}\epsilon_{stat.,s^{n}_{l}}^{post.}\big{)},$ (23)
where $\mathsf{M}\big{(}\epsilon_{bias,s^{n}_{l}}\big{)}$, $\mathsf{M}\big{(}\epsilon_{stat.,s^{n}_{l}}\big{)}$ and $\mathsf{M}\big{(}\epsilon_{stat.,s^{n}_{l}}^{post.}\big{)}$ are respectively upper bounds for the bias error of the mitigated probability $\epsilon_{bias,s^{n}_{l}}$, the statistical error of the mitigated probability $\epsilon_{stat.,s^{n}_{l}}$, and the statistical error of the postselected probability $\epsilon_{stat.,s^{n}_{l}}^{post.}$. More explicitly, in the appendix 18 we show the following
Theorem 10.
The condition:
$O\bigg{(}\sqrt{\frac{{m\choose n}}{{n\choose k}(1-\eta)^{n-k}\eta^{k}N_{tot}}}\bigg{)}+\bigg{(}{m-n+k\choose k}-1\bigg{)}O\bigg{(}\frac{1}{\mathsf{poly}(m)}\bigg{)}\leq O\Bigg{(}\sqrt{\frac{{m\choose n}}{(1-\eta)^{n}N_{tot}}}\Bigg{)},$ (24)
defines a sampling regime (a range of values of $N_{tot}$) where the sum of the worst-case statistical error and bias error of linear solving recycling mitigation is less than the worst-case statistical error of postselection.
Note that the big O notation here can be interpreted as multiplying by a prefactor, which can be computed by choosing a confidence for the concentration inequalities (as will be seen in section 9). This allows for the direct computation of the bounds for a specific confidence. One can intuitively understand Thm. 10 as follows. For values of loss $\eta>0.5$, it is expected that the statistical error of recycling mitigation $O\bigg{(}\sqrt{\frac{{m\choose n}}{{n\choose k}(1-\eta)^{n-k}\eta^{k}N_{tot}}}\bigg{)}$ is in general smaller than that of postselection $O\bigg{(}\sqrt{\frac{{m\choose n}}{(1-\eta)^{n}N_{tot}}}\bigg{)}$. Thus, the sampling regime where recycling outperforms postselection is typically determined by the bias error $2\big{(}{m-n+k\choose k}-1\big{)}O\big{(}\frac{1}{\text{poly}(m)}\big{)}\approx O\big{(}\frac{1}{\mathsf{poly}(m)}\big{)}$, for small values of $k$ independent of $n$. This means that recycling is expected to outperform postselection for estimating output probabilities to up to inverse polynomial errors, with high confidence. Furthermore, Thm. 9 shows that, for many linear optical circuits, the outperformance is up to inverse exponentially small errors.
We now introduce the notion of dependency. This quantifies the correlation of the interference terms with the ideal probability within the recycled probabilities. A positive correlation means that the signal for the ideal probability is greater than (${{m-n+k}\choose k}^{-1}$, which can be used to improve the performance of linear solving. Each recycled probability can be rewritten to include a dependency term $d_{k}(s_{l}^{n})$ quantifying this correlation, by reformulating the interference term as a linear function of $p(s_{l}^{n})$ and $p_{unif}$. The expression
@@
defines the dependency term $d_{k}(s^{n}_{l})$ of each recycled probability, and $k\leq n-1$. An average dependency term over the distribution, denoted $d_{k}$, may be calculated from the recycled probabilities (see eqn. 18.2.1 and appendix 18). Each recycled probability constructed from the $n-k$ output photon statistics may then be expressed in the form
$\begin{split}\tilde{p}_{R}^{k}(s^{n}_{l})&=\frac{p(s^{n}_{l})}{{m-n+k\choose k}}+\frac{N^{\prime}_{k}}{N_{k}}\big{(}\big{(}(1-d_{k})p_{unif}+d_{k}p(s^{n}_{l})+\epsilon_{bias,s^{n}_{l}}\big{)}+\epsilon_{stat.,s^{n}_{l}},\\ \end{split}$ (26)
@@
input : Recycled probability estimator $\tilde{p}^{k}_{R}(s^{n}_{l})$, uniform probability $p_{unif}$, an estimator for the absolute average deviation for the $n$ output photon distribution $\tilde{D}_{0}$, and an estimator for the absolute average deviation for the $n-k$–photon recycled distribution $\tilde{D}_{k}$ (eqn. (29)).
1  Initialise variables for the mitigated output $p_{mit}(s^{n}_{l})\leftarrow 0$ and the average dependency term $d_{k}\leftarrow 0$.
2 Update the average dependency term variable $d_{k}\leftarrow\frac{1}{{{m-n+k}\choose k}-1}\bigg{(}\frac{{{m-n+k}\choose k}\tilde{D}_{k}}{\tilde{D}_{0}}-\frac{1}{{{m-n+k}\choose k}}\bigg{)}$.
3 Update the mitigated output variable $p_{mit}(s^{n}_{l})\leftarrow\Bigg{|}\frac{\tilde{p}_{R}^{k}(s^{n}_{l})+\frac{N^{\prime}_{k}}{N_{k}}\big{(}(-1+d_{k})p_{unif}\big{)}}{{{m-n+k\choose k}}^{-1}+\frac{N^{\prime}_{k}}{N_{k}}d_{k}}\Bigg{|}$
@@
In appendix 18, we show the following.
Theorem 11.
The condition
$O\bigg{(}\sqrt{\frac{{m\choose n}}{{n\choose k}(1-\eta)^{n-k}\eta^{k}N_{tot}}}\bigg{)}+2\bigg{(}{m-n+k\choose k}-1\bigg{)}O\bigg{(}\frac{1}{\mathsf{poly}(m)}\bigg{)}\leq O\Bigg{(}\sqrt{\frac{{m\choose n}}{(1-\eta)^{n}N_{tot}}}\Bigg{)},$ (27)
defines a sampling regime where the sum of the worst-case statistical error and bias error of linear solving with dependency recycling mitigation is less than the worst-case statistical error of postselection.
The performance guarantees for linear solving with and without dependency are too similar to be used to draw inferences on their relative performance. We do note, however, that in numerical simulation linear solving with dependency reliably outperforms linear solving without dependency (see Fig. 3 (a) and (b)).
8.2 Extrapolation
@@
the magnitude of the ideal probability signal within the recycled probabilities is proportional to (${m-n+k\choose k}^{-1}$. As $k$ increases, the ideal signal magnitudes decrease as the recycled distributions converge towards uniform. The rate of decay of the ideal signal with $k$ can be computed and then used to extrapolate mitigated outputs. We present two variations of extrapolation in which different types of dependence of the ideal probability signal on number of lost photons $k$ are considered. A linear dependence is used for a linear extrapolation method, and an exponential dependence for an exponential extrapolation method. Both the linear and exponential extrapolation methods involve two iterations of optimisation. The first iteration is to compute an average decay parameter using the set of average absolute deviations of the recycled distributions. Where, for the set of recycled probabilities $\{p^{k}_{R}(s^{n}_{l})\}_{l}$ constructed from $k$ photon statistics, the average absolute deviation is defined as
$\begin{split}D_{k}&:={m\choose n}^{-1}\sum_{l}\absolutevalue{{p}_{R}^{k}(s^{n}_{l})-p_{unif}}.\\ \end{split}$ (29)
@@
where $\tilde{D}_{0}$ is the average absolute deviation of the $n$–photon distribution from uniform computed from the statistics corresponding to postselecting on detecting all $n$ photons, and $g_{\text{avg}}$ is the optimal global linear decay parameter. The set $\{k,\tilde{D}_{k}\}_{k=1}^{K}$ is used as the data set to compute $g_{\text{avg}}$. Where $\{k,\tilde{D}_{k}\}_{k=1}^{K}$ is set of average absolute deviations from uniform for the different distributions, for $K\leq n$ and $\tilde{D}_{k}$ is the estimated (from statistics where $k$ photons were lost) absolute average deviation for the $n-k$–photon recycled distribution. After the optimal decay parameter is identified, another iteration of least squares is performed with updated model functions this time to generate the mitigated values. The data set used in this step is $\{k,|\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}|\}^{K}_{k=1}$. For the second iteration of linear least squares, each output bit string is assigned a linear model function of the form
$f_{s_{n}}(x,\alpha_{s^{n}_{l}})=\text{sgn}(p_{unif}-y_{1})g_{\text{avg}}x+\alpha_{s^{n}_{l}}.$ (32)
@@
For each output bit string $s_{n}$ an optimal $\alpha_{s^{n}_{l}}$ is computed, generating the set $\{\alpha_{s^{n}_{l}}\}_{l}$, and the set of mitigated outputs is then $\{\alpha_{s^{n}_{l}}+p_{unif}\}_{l}$. We now provide pseudocode for applying the linear extrapolation method to generate mitigated outputs.
@@
input : The number of data points $n_{d}\in\{n_{d}\in\mathbb{Z}^{+}|n_{d}<n\}$ to be used in both iterations of least squares, the data set $\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ used to compute the gradient parameter $g_{\text{avg}}$ in the first iteration of least squares, $\tilde{D}_{0}$, and the data set $\{k,|\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}|\}_{k=1}^{n_{d}}$ used to compute the mitigated output in the second iteration of least squares.
@@
1  Initialise an average decay parameter variable $\tilde{g}_{\text{avg}}\leftarrow 0$, a prefactor variable $\alpha_{s^{n}_{l}}\leftarrow 0$, and a mitigated output variable $p_{mit}(s^{n}_{l})\leftarrow 0$.
@@
2 Use least squares method with model function $f(x_{i},g_{\text{avg}})=-\tilde{g}_{\text{avg}}x_{i}+\tilde{D}_{0}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ to compute the value of the average decay parameter (slope), and assign this to variable $\tilde{g}_{\text{avg}}$.
@@
3 Use least squares method with model function $f_{s^{n}_{l}}(x_{i},\alpha_{s^{n}_{l}})=\text{sgn}(p_{unif}-y_{1})\tilde{g}_{\text{avg}}x_{i}+\alpha_{s^{n}_{l}}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,|\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}|\}_{k=1}^{n_{d}}$ to compute the value of the y-axis intercept, and assign this to variable $\alpha_{s^{n}_{l}}$.
4 Update mitigated output variable as $p_{mit}(s^{n}_{l})\leftarrow p_{unif}+\alpha_{s^{n}_{l}}$.
@@
The analytic condition for linear extrapolation to outperform postselection is now stated.
Theorem 12.
The condition:
$\begin{split}\frac{n_{d}+1}{2}\bigg{(}O\bigg{(}\bigg{(}\frac{1}{(1-\eta)^{n}N_{tot}}&\bigg{)}^{1/4}\bigg{)}+O\Big{(}\frac{1}{\text{poly}(m)}\Big{)}\bigg{)}\\ &+O\bigg{(}\frac{n}{m-n+1}\sqrt{\frac{{m\choose n}}{n(1-\eta)^{n-1}\eta N_{tot}}}\bigg{)}\leq O\bigg{(}\sqrt{\frac{{m\choose n}}{(1-\eta)^{n}N_{tot}}}\bigg{)},\end{split}$ (33)
defines a sampling regime where the sum of the worst-case statistical error and bias error of linear extrapolation recycling mitigation using the least squares method is less than the worst-case statistical error of postselection.
Thm. 12 is proven in appendix 19. The left hand side of the above expression is a sum of the upper bounds of the errors for performing linear extrapolation on the recycled probabilities. Using left-to-right ordering, the first term and third terms are due to statistical error and the second term is due to bias error. While the single term on the right side of the inequality is due to the statistical error of the postselected output.
@@
where $\alpha_{\text{avg}}$ is the optimal global exponential decay parameter. Where again the set $\{k,\tilde{D}_{k}\}_{k=1}^{K}$ is used as the data, this time to compute the optimal value of $\alpha_{\text{avg}}$. And then each output bit string for which a mitigated probability is to be generated is assigned a model function of the form
@@
input : The number of data points $n_{d}\in\{n_{d}\in\mathbb{Z}^{+}|n_{d}<n\}$ to be used in both iterations of least squares, the set $\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ used to compute the gradient parameter $g_{\text{avg}}$ in the first iteration of least squares, $\tilde{D}_{0}$, and the data set $\{k,|\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}|\}_{k=1}^{n_{d}}$ used to compute the prefactor value in the second iteration of least squares.
1  Initialise an average decay parameter variable $\alpha_{avg}\leftarrow 0$, a prefactor variable $\Lambda_{s^{n}_{l}}\leftarrow 0$, and a mitigated output variable $p_{mit}(s^{n}_{l})\leftarrow 0$.
2 Use least squares method with model function $f(x,\alpha_{\text{avg}})=\tilde{D}_{0}e^{-\alpha_{\text{avg}}x}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ to compute the value of the average decay parameter and assign this to variable ${\alpha}_{\text{avg}}$.
3 Use least squares method with model function $f_{s_{n}}(x,\Lambda_{s^{n}_{l}})=\Lambda_{s^{n}_{l}}e^{-\alpha_{\text{avg}}x}+p_{unif}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,|\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}|\}_{k=1}^{n_{d}}$ to compute the value of the prefactor and assign this to variable $\Lambda_{s^{n}_{l}}$.
4 Update mitigated output variable as $p_{mit}(s^{n}_{l})\leftarrow p_{unif}+\Lambda_{s^{n}_{l}}$.

### v3

@@ 4.4.1 Linear solving
1ex Initialise variable $p_{\text{miti}}(s^{n}_{l})\leftarrow 0$ for the mitigated output to be computed.
Update mitigated output variable as $p_{\text{miti}}(s^{n}_{l})\leftarrow{{m-n+k\choose k}}\big|\tilde{p}_{R}^{k}(s^{n}_{l})-\frac{N^{\prime}_{k}}{N_{k}}p_{unif}\big|$.
@@
There exists a regime of recycling mitigation usefulness, where the combined bias and the statistical errors present in the mitigated probabilities are lower than the statistical errors of the postselected distribution. Indeed, from Equation 28, it can be seen that the bias error of linear solving is upper bounded by
$M_{bias}\in O\bigg({m-n+k\choose k}\bigg)\epsilon_{bias,s^{n}_{l}},$
we formally prove this in Appendix 12. Similarly, and as stated in Section 3, we show in Appendix 12 that statistical errors for a recycled probability constructed from nkn-k photon statistics in linear solving is upper bounded by $M_{stat,k}\in O\bigg(\sqrt{\frac{{m\choose n}}{{n\choose k}(1-\eta)^{n-k}\eta^{k}N_{tot}}}\bigg)$, the $m\choose n$} can be replaced with $N_{s}$, if one is interested in mitigating a subset $N_{s}$ of probabilities. With these upper bounds in hand, one can find a condition on the number of samples up to which recycling mitigation outperforms postselection by solving for $N_{tot}$ in the following inequality
$M_{bias}+M_{stat,k}\leq M_{stat,k=0}.$
More details on this can be found in Appendix 12.
We now introduce the notion of dependency. This quantifies the correlation of the interference terms with the ideal probability within the recycled probabilities. A positive correlation means that the signal for the ideal probability is greater than ${{m-n+k}\choose k}^{-1}$, which can be used to improve the performance of linear solving. Each recycled probability can be rewritten to include a dependency term $d_{k}(s_{l}^{n})$ quantifying this correlation, by reformulating the interference term as a linear function of $p(s_{l}^{n})$ and $p_{unif}$. The expression
@@
defines the dependency term $d_{k}(s^{n}_{l})$ of each recycled probability, and $k\leq n-1$. An average dependency term over the distribution, denoted $d_{k}$, may be calculated from the recycled probabilities (see eqn. 12.3.1 and appendix 12). Each recycled probability constructed from the nkn-k output photon statistics may then be expressed in the form
[…]
@@
We note that although the upper-bounds derived in Appendix 12 on the statistical and bias errors of linear solving with dependency are similar to those for linear solving without dependency, nevertheless numerical simulations show that linear solving with dependency reliably outperforms linear solving without dependency (see Fig. 3 (a) and (b)).
input : Recycled probability estimator $\tilde{p}^{k}_{R}(s^{n}_{l})$, uniform probability $p_{unif}$, an estimator for the absolute average deviation for the $n$ output photon distribution $\tilde{D}_{0}$, and an estimator for the absolute average deviation for the nkn-k–photon recycled distribution $\tilde{D}_{k}$ (eqn. (32)).
1ex Initialise variables for the mitigated output $p_{mit}(s^{n}_{l})\leftarrow 0$ and the average dependency term $d_{k}\leftarrow 0$.
Update the average dependency term variable $d_{k}\leftarrow\frac{1}{{{m-n+k}\choose k}-1}\bigg(\frac{{{m-n+k}\choose k}\tilde{D}_{k}}{\tilde{D}_{0}}-\frac{1}{{{m-n+k}\choose k}}\bigg)$.
Update the mitigated output variable $p_{mit}(s^{n}_{l})\leftarrow\Bigg|\frac{\tilde{p}_{R}^{k}(s^{n}_{l})+\frac{N^{\prime}_{k}}{N_{k}}\big((-1+d_{k})p_{unif}\big)}{{{m-n+k\choose k}}^{-1}+\frac{N^{\prime}_{k}}{N_{k}}d_{k}}\Bigg|$
@@
4.4.2 Extrapolation
@@
the magnitude of the ideal probability signal within the recycled probabilities is proportional to ${m-n+k\choose k}^{-1}$. As $k$ increases, the ideal signal magnitudes decrease as the recycled distributions converge towards uniform. The rate of decay of the ideal signal with $k$ can be computed and then used to extrapolate mitigated outputs. We present two variations of extrapolation in which different types of dependence of the ideal probability signal on number of lost photons $k$ are considered. A linear dependence is used for a linear extrapolation method, and an exponential dependence for an exponential extrapolation method. These two functions were considered primarily for heuristic reasons as seeming to reflect the decay behaviour observed in the recycled distributions. Both the linear and exponential extrapolation methods involve two iterations of optimisation. The first iteration computes an average decay parameter using the set of average absolute deviations of the recycled distributions. Where, for the set of recycled probabilities $\{p^{k}_{R}(s^{n}_{l})\}_{l}$ constructed from $k$ photon statistics, the average absolute deviation is defined as
[…]
@@
where $\tilde{D}_{0}$ is the average absolute deviation of the $n$-photon distribution from uniform computed from the statistics corresponding to postselecting on detecting all $n$ photons, and $g_{\text{avg}}$ is the optimal global linear decay parameter. The set $\{k,\tilde{D}_{k}\}_{k=1}^{K}$ is used as the data set to compute $g_{\text{avg}}$. Where $\{\tilde{D}_{k}\}_{k=1}^{K}$ is the set of average absolute deviations from uniform for the different distributions, for $K\leq n$, and $\tilde{D}_{k}$ is the estimated (from statistics where $k$ photons were lost) absolute average deviation for the nkn-k–photon recycled distribution. After the optimal decay parameter is identified, another iteration of least squares is performed with updated model functions this time to generate the mitigated values. The data set used in this step is $\{k,\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}\}^{K}_{k=1}$. For the second iteration of linear least squares, each output bit string is assigned a linear model function of the form
$f_{s_{n}}(x,\alpha_{s^{n}_{l}})=\text{sgn}(p_{unif}-\tilde{p}^{k=1}_{R}(s^{n}_{l}))g_{\text{avg}}x+\alpha_{s^{n}_{l}}.$ (35)
For each output bit string $s_{n}$ an optimal $\alpha_{s^{n}_{l}}$ is computed, generating the set $\{\alpha_{s^{n}_{l}}\}_{l}$, and the set of mitigated outputs is then $\{\alpha_{s^{n}_{l}}+p_{unif}\}_{l}$. In Appendix 13, error bounds are derived for the use of linear extrapolation to perform recycling mitigation, these indicate the existence of a regime where the mitigation outperforms postselection. As with the analytical linear solving error bounds, the linear extrapolation error bounds allow the computation of an estimate of the number of samples up to which recycling mitigation with linear extrapolation outperforms postselection by solving for $N_{tot}$ in the following inequality
$M_{bias}+M_{stat,k}\leq M_{stat,k=0}.$
@@
We now provide pseudocode for applying the linear extrapolation method to generate mitigated outputs.
@@
input : The number of data points $n_{d}\in\{n_{d}\in\mathbb{Z}^{+}|n_{d}<n\}$ to be used in both iterations of least squares, the data set $\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ used to compute the gradient parameter $g_{\text{avg}}$ in the first iteration of least squares, $\tilde{D}_{0}$, and the data set $\{k,\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}\}_{k=1}^{n_{d}}$ used to compute the mitigated output in the second iteration of least squares.
@@
1ex Initialise an average decay parameter variable $\tilde{g}_{\text{avg}}\leftarrow 0$, a prefactor variable $\alpha_{s^{n}_{l}}\leftarrow 0$, and a mitigated output variable $p_{mit}(s^{n}_{l})\leftarrow 0$.
@@
Use least squares method with model function $f(x_{i},g_{\text{avg}})=-\tilde{g}_{\text{avg}}x_{i}+\tilde{D}_{0}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ to compute the value of the average decay parameter (slope), and assign this to variable $\tilde{g}_{\text{avg}}$.
@@
Use least squares method with model function $f_{s^{n}_{l}}(x_{i},\alpha_{s^{n}_{l}})=\text{sgn}(p_{unif}-\tilde{p}_{R}^{k=1}(s^{n}_{l}))\tilde{g}_{\text{avg}}x_{i}+\alpha_{s^{n}_{l}}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}\}_{k=1}^{n_{d}}$ to compute the value of the y-axis intercept, and assign this to variable $\alpha_{s^{n}_{l}}$.
Update mitigated output variable as $p_{mit}(s^{n}_{l})\leftarrow p_{unif}+\alpha_{s^{n}_{l}}$.
@@
where $\alpha_{\text{avg}}$ is the optimal global exponential decay parameter  [^3] Note that, in order to satisfy $f(n,\alpha_{\mathsf{avg}})=0$, the fitting function should be $f(x,\alpha_{\mathsf{avg}})=\tilde{D}_{0}e^{-\alpha_{\text{avg}}x}+\beta$ with $\beta<0$. However, as we are intersted in the decay of $f(.)$ with increasing $k$, and we look at (far from $n$) values of $k\in O(1)$ in the fitting, we can exclude $\beta$ in the above expression. . Where again the set $\{k,\tilde{D}_{k}\}_{k=1}^{K}$ is used as the data, this time to compute the optimal value of $\alpha_{\text{avg}}$. And then each output bit string for which a mitigated probability is to be generated is assigned a model function of the form
@@
input : The number of data points $n_{d}\in\{n_{d}\in\mathbb{Z}^{+}|n_{d}<n\}$ to be used in both iterations of least squares, the set $\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ used to compute the gradient parameter $g_{\text{avg}}$ in the first iteration of least squares, $\tilde{D}_{0}$, and the data set $\{k,\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}\}_{k=1}^{n_{d}}$ used to compute the prefactor value in the second iteration of least squares.
1ex Initialise an average decay parameter variable $\alpha_{avg}\leftarrow 0$, a prefactor variable $\Lambda_{s^{n}_{l}}\leftarrow 0$, and a mitigated output variable $p_{mit}(s^{n}_{l})\leftarrow 0$.
Use least squares method with model function $f(x,\alpha_{\text{avg}})=\tilde{D}_{0}e^{-\alpha_{\text{avg}}x}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,\tilde{D}_{k}\}_{k=1}^{n_{d}}$ to compute the value of the average decay parameter and assign this to variable ${\alpha}_{\text{avg}}$.
Use least squares method with model function $f_{s_{n}}(x,\Lambda_{s^{n}_{l}})=\Lambda_{s^{n}_{l}}e^{-\alpha_{\text{avg}}x}+p_{unif}$ and data set $\{x_{i},y_{i}\}_{i=1}^{n_{d}}:=\{k,\tilde{p}^{k}_{R}(s^{n}_{l})-p_{unif}\}_{k=1}^{n_{d}}$ to compute the value of the prefactor and assign this to variable $\Lambda_{s^{n}_{l}}$.
Update mitigated output variable as $p_{mit}(s^{n}_{l})\leftarrow p_{unif}+\Lambda_{s^{n}_{l}}$.